# Conveying Emotions to Robots through Touch and Sound

### v1

@@ Abstract
Human emotions can be conveyed through nuanced touch gestures. However, there is a lack of understanding of how consistently emotions can be conveyed to robots through touch. This study explores the consistency of touch-based emotional expression toward a robot by integrating tactile and auditory sensory reading of affective haptic expressions. We developed a piezoresistive pressure sensor and used a microphone to mimic touch and sound channels, respectively. In a study with 28 participants, each conveyed 10 emotions to a robot using spontaneous touch gestures. Our findings reveal a statistically significant consistency in emotion expression among participants. However, some emotions obtained low intraclass correlation values. Additionally, certain emotions with similar levels of arousal or valence did not exhibit significant differences in the way they were conveyed. We subsequently constructed a multi-modal integrating touch and audio features to decode the 10 emotions. A support vector machine (SVM) model demonstrated the highest accuracy, achieving 40% for 10 classes, with “Attention” being the most accurately conveyed emotion at a balanced accuracy of 87.65 %.
Keywords:
Tactile interaction affective computing emotion decoding multi-modal1 Introduction
Touch is essential to human social interactions, serving as a vital medium for social exchange. Through touch, individuals form strong bonds and cooperative relationships, navigate social hierarchies, provide comfort and reassurance, and convey romantic interest [1]. A simple touch can communicate comfort, reassurance, empathy, or even urgency without a single word being spoken. The physical contact involved in touch encompasses a range of parameters, including pressure, duration, frequency, and location, all of which contribute to the pro-social emotional message being conveyed [1]. For instance, a light, brief touch on the shoulder may express sympathy, while a firm, prolonged grip can indicate support and solidarity [2].
Advancements in robotics and sensor technology have enabled more complex human-robot interactions. These sensors help us build robots with tactile sensitivity, allowing them to detect and respond to human emotions through touch. Tactile sensors capture nuances like pressure and texture, vital for emotion recognition [3], while auditory feedback from touch gestures enhances emotion decoding accuracy [4][5]. Several studies have explored how distinct emotions can be conveyed through various touch gestures [6]. For example, Yohanan and MacLean developed the Haptic Creature to detect touch and movement, investigating which gestures from a predefined set participants believed were appropriate for expressing nine specific emotions [3]. Previous research has attempted to decode emotions using video recordings of users touching robots but has often neglected crucial tactile elements such as the applied force [7]. Some studies instructed participants to express specific emotions through touch to see if subjective feelings could be communicated via simple vibrotactile feedback derived from touch sounds [4]. These studies primarily focused on either tactile or auditory channels in isolation, and focused on the correlations between touch gestures and emotions.
In this work, we collected data from touch and sound channels for the robot to mimic human touch perception. Specifically, we designed a tactile sensor to collect data from touch channels and additionally collect auditory signals from touch sounds. We explored the consistency of emotion expression based on touch and sound features using multi-modal methods, including sound and touch channels. We focus on the following research questions: (1) Are people inter-consistent in the way they express an emotion through touch? (2) Which emotions are distinguishable when expressed through touch? (3) Can multi-modal be used to decode different emotions?
2 Materials & Methods
2.1 Equipments
@@
In this experiment, participants were tasked with conveying emotions to the Pepper robot, see Fig. 1. The participants were presented with definitions of the emotions to ensure they agreed with and understood the given definitions. The data collection involved the use of a custom 5-by-5 tactile sensor and a microphone to capture gesture data. Fig. 1(a) shows how the sensors are mounted on the Pepper robot. The following sections will detail the equipment used, the setup process, and the specific configurations implemented for data acquisition.
2.1.1 Tactile sensor
@@
The tactile sensor is a 5-by-5 piezoresistive pressure grid, similar to the Smart Textile we developed in [8]. As shown in Fig 2(a), the sensor consists of a top and bottom electrode pattern with a piezoresistive Velostat™ sheet in between. The Velostat™ sheet was roughened by pressing it between two pieces of 80 grit sanding paper: this ensures little contact between the Velostat™ and the electrodes when no pressure is applied, and thus a small nominal sensor reading. The electrodes are made by screenprinting DuPont™ PE874™ conductive ink onto a Bemis™ 3914 thermoplastic polyurethane (TPU) sheet. For each electrode, the printed side is facing the Velostat™ sheet. To interface the printed TPU with a printed circuit board (PCB) for readout, flex PCBs are glued to the printed TPU sheet. This is achieved by adding extra conductive ink on the traces and Gorilla® Super Glue in between. Finally, extra TPU is used to cover the remaining exposed traces, to prevent oxidisation. This was omitted from Fig. 2 for clarity of the graphic. The entire structure is fused together using a Hotronix® Air Fusion IQ® heat press at and 1.4 bar for .
@@
The readout PCB, shown in Fig.2(b), uses a NINA-B306 microcontroller unit (MCU), powered by an external lithium-ion polymer (LiPo) battery. Flexible printed cable (FPC) connectors are used to interface with the flex PCBs. The MCU communicates the sensor data wirelessly over Bluetooth Low Energy (BLE) at an average rate of 45 Hz. Custom Arduino board files were developed for the PCB so that it can be programmed using the Arduino IDE. We provide a guide for creating such a custom Arduino board in our GitHub repository111https://github.com/RemkoPr/airo-nrf52840-boards.
@@
The transduction mechanism of the sensor is as follows. When pressure is applied to the sensor, the resistance of the Velostat sheet decreases. Hence, as shown in Fig.2(c), the sensor can be modeled as a grid of variable resistors. Each of the traces of the top electrode is connected to a digital pin D0..5 of the MCU. The traces of the bottom electrode run into a demultiplexer (DEMUX), of which the output is connected to an analog pin A0 of the MCU. This pin has a pulldown resistor. Such a low pulldown value biases the sensor as a sensitive touch sensor, rather than a pressure sensor with large dynamic range. To read a single taxel, one of the digital pins is set high (3.3 V), the others to a high impedance (Hi-Z) state. The DEMUX then routes one of the bottom traces to the A0 pin. The voltage at this pin will depend on both the pulldown resistor as well as the resistance of the Velostat™ squeezed between the top trace that is driven high, and the bottom trace that is selected by the DEMUX:
2.1.2 Microphone
@@
2.1.3 Definitions of emotions
Affective states can be represented along dimensions of arousal and valence, as described by Russell’s model [9]. In our study, we selected nine distinct emotions from different arousal and valence zones, plus one neutral emotion, “anger”, “fear”, and “disgust” fall within the high arousal and negative valence quadrant. “Happiness” and “surprise” are in the high arousal and positive valence quadrant. “sadness” and “confusion” belong to the low arousal and negative valence quadrant, while “comfort” and “calm” are from the low arousal and positive valence quadrant. “attention” is categorized as a neutral emotion. Before data collection, participants were provided with definitions of these ten emotions to ensure they understood and agreed with the interpretations. Details of each emotion definition refer to the Oxford dictionary and can be found at this GitHub repository222https://github.com/qiaoqiao2323/Convey-emotion.
2.2 Experimental design
@@
To explore emotion decoding via touch-based expression and measure the consistency with which emotions and gestures are expressed across different individuals, we set up a data collection experiment. We recruited all the participants from a similar cultural background. Specifically, the group comprises 28 Chinese participants, 10 identified as female, 18 as male, and their average age is years. The data collection and study adhered to the ethics procedures of Ghent University, and participants gave informed consent. Participants were asked to convey each of the 10 emotions to the robot, using a gesture of their own choosing to express each emotion, with each gesture being recorded for a maximum of 10 seconds, as referred in [4]. This was repeated 3 times, so for each emotion, 3 recordings per participant were made. Participants were given time between each emotion to prepare and think about their next performance. The subsequent data collection only commenced once the participants indicated they were ready. Participants completed three rounds, with each subsequent round commencing only after all emotions or gestures had been performed, ensuring a greater variety in the data collected. After gesture data collection, we also gathered subjective feedback by asking each participant which emotions were more challenging to express.
3 Data analysis
3.1 Dataset
@@
The dataset consists of tactile sensor and microphone recordings collected to study the conveyance of emotions through touch and audio. It includes data from 28 participants, each performing three rounds of interactions. For each of the 10 emotions, there are a total of 84 samples (28 participants for 3 rounds). Each interaction lasted 10 seconds. Tactile sensor data were sampled at a rate of 45 Hz, providing continuous feedback during the interactions. Simultaneously, audio data were recorded in 10-second WAV files.
3.2 Audio Features
@@ 3.3 Tactile Features
Median Force, representing the middle value of force applied, IQR Force measures the dispersion of the middle 50% of force values. Contact Area is crucial in identifying the nature of the touch, with larger areas potentially indicating comforting emotions, while smaller areas might suggest precision or nervousness. The Rate of Pressure Change captures how quickly the pressure varies, with high rates linked to emotions like excitement. Pressure Std reflects the fluctuation in pressure, and Numtouches quantifies the frequency of touch events within 10 seconds. Max Touch Duration, Min Touch Duration, and Mean Touch Duration describe the temporal aspects of touch, indicating how long touches are sustained, which can be correlated with different emotional expressions.
3.4 Subjective feedback
We gathered subjective feedback by asking participants which emotions were more challenging to express (multiple choice). Surprise and confusion ranked the highest, with 64.3% and 46.4% of participants choosing these emotions, respectively, out of the 10 emotions evaluated.
3.5 Consistency analysis

### v2

# Touch and Tell: Multimodal Decoding of Human Emotions and Social Gestures for Robots

@@ Abstract
Human emotions are complex and can be conveyed through nuanced touch gestures. Previous research has primarily focused on how humans recognize emotions through touch or on identifying key features of emotional expression for robots. However, there is a gap in understanding how reliably these emotions and gestures can be communicated to robots via touch and interpreted using data-driven methods. This study investigates the consistency and distinguishability of emotional and gestural expressions through touch and sound. To this end, we integrated a custom piezoresistive pressure sensor as well as a microphone on a social robot. Twenty-eight participants first conveyed ten different emotions to the robot using spontaneous touch gestures, then they performed six predefined social touch gestures. Our findings reveal statistically significant consistency in both emotion and gesture expression among participants. However, some emotions exhibited low intraclass correlation values, and certain emotions with similar levels of arousal or valence did not show significant differences in their conveyance. To investigate emotion and social gesture decoding within affective human-robot tactile interaction, we developed single-modality models and multimodal models integrating tactile and auditory features. A support vector machine (SVM) model trained on multimodal features achieved the highest accuracy for classifying ten emotions, reaching 40 %.For gesture classification, a Convolutional Neural Network - Long Short-Term Memory Network (CNN-LSTM) achieved 90.74 % accuracy. Our results demonstrate that even though the unimodal models have the potential to decode emotions and touch gestures, the multimodal integration of touch and sound significantly outperforms unimodal approaches, enhancing the decoding of both emotions and gestures.
Index Terms:
Tactile interaction, affective computing, sound communication, multimodal, human-robot interaction, emotion classification, gesture classification.I Introduction
Understanding human emotions is a fundamental goal in the development of socially interactive robots. As robots become increasingly integrated into domains such as healthcare, therapy, customer service, and companionship, their ability to perceive and appropriately respond to human affect is becoming essential. Affective computing, which focuses on systems that recognize and process emotional signals, plays a critical role in enabling more natural and compelling human-robot interaction [1]. Emotion recognition has been widely explored through modalities such as facial expressions, speech, and body gestures [2]. Previous research has shown that intelligent virtual agents can be endowed with empathic behaviours to serve in roles such as virtual therapists [3]. However, affect recognition through touch remains relatively underexplored, despite its central importance in human social communication.
Touch serves as a powerful and nuanced channel for emotional and social exchange. It plays a vital role in forming interpersonal bonds, navigating social roles, offering reassurance, and conveying empathy or urgency [4]. The communicative power of touch lies in its parameters—pressure, duration, frequency, and location—which together shape the emotional message being conveyed [4, 5]. For instance, a brief pat may express comfort, whereas a firm, prolonged grip may signal support or solidarity.
Recent studies have increasingly highlighted the emotional and communicative value of touch in interactive systems. For example, emotional expressions of virtual agents have been shown to influence how humans physically interact with them in VR settings [6], and tele-touch devices have enabled the remote transmission of affect [7]. Affective cues have also been inferred from everyday smartphone touch gestures, such as taps and swipes, to assess stress levels [8]. Furthermore, agents combining touch with facial expressions and speech can convey empathy more effectively, though in some contexts, touch alone may be the most influential cue [9]. Experimental research using robotic arms has shown that specific touch features (e.g., force, speed, amplitude) shape perceptions of arousal and valence, and can even override conflicting emotional cues from other modalities [10]. Physiological studies further reveal that social touch elicits distinct autonomic responses compared to self- or object-touch, indicating its unique role in emotional communication [11]. Moreover, affective touch delivered through a robot has been shown to reduce users’ heart and respiration rates and increase their sense of calm and happiness, reinforcing its therapeutic potential [12].
In addition to emotional expression, researchers have studied the classification of social touch gestures, recognizing their capacity to communicate intent, build intimacy, and support remote interaction [13]. People can express meaning through structured haptic codes [14], and touch-based gesture decoding has emerged as a growing field. The Social Touch Gesture Challenge 2015, for instance, provided annotated tactile data on various social gestures performed on pressure-sensitive surfaces of different textures and shapes [15]. This lays the groundwork for building systems capable of decoding the social and emotional meaning behind touch, a necessary step toward enabling robots to respond to human touch in a contextually appropriate and socially meaningful way.
These findings collectively underscore the importance of decoding the emotional and social meaning of touch gestures, both as an independent modality and as part of multimodal interaction, to develop socially aware and emotionally intelligent robots.
I-A Robot touch sensing capabilities
Advancements in robotics and sensor technology have enabled more complex human-robot interactions. These sensors help us build robots with tactile sensitivity, allowing them to detect and respond to human emotions through touch. Tactile sensors capture nuances like pressure and texture, vital for emotion recognition [16], while auditory feedback from touch gestures enhances emotion decoding accuracy [17]. One approach for creating robot touch sensing capabilities involves covering a robot with artificial skin that simulates the human touch sensory system[18]. Alternatively, robots can be embedded with sensors to measure temperature, proximity, and pressure[19].
Previous research designed soft tactile sensors, which are equipped on the iCub robot and successfully used for shape exploration [20, 21]. Naya et al. built an artificial skin by covering the robot with a gridded pressure-sensitive ink sheet [22]. Using features such as absolute pressure and temporal pressure differences, they distinguished five touch gestures: “Pat”, “Scratch”, “Slap”, “Stroke”, and “Tickle”. Silvera et al. further developed the concept of artificial skin using electrical impedance tomography, which allowed larger areas of a robot to be covered and enabled the extraction of information such as location, duration, displacement, and intensity of touch [23]. In addition to academic developments in tactile sensing, the industry has also explored ways to equip robots with sensing capabilities. For example, solutions like AIRSKIN [24] and Dobot SafeSkin [25] aim to enable safe and responsive physical interaction.
Simulating touch perception capabilities using sensors has already been extensively applied in social robotics. For example, the Huggable is a robotic companion designed in the form of a teddy bear, equipped with a combination of temperature, electric field, and force sensors to facilitate affective haptic communication between individuals [26]. The Haptic Creature is a furry, lap-sized social robot embedded with pressure sensors throughout its body, enabling it to communicate with the world through touch [27]. Paro, a social robot modelled after a seal, is equipped with sensors that allow it to respond to touch and distinguish between three different gestures: petting, stroking, and hitting [28].
I-B Related work
I-B1 Conveying emotions
Several studies have explored how distinct emotions can be conveyed through various touch gestures [29, 30, 10]. For example, Yohanan and MacLean developed the Haptic Creature to detect touch and movement, investigating which gestures from a predefined set participants believed were appropriate for expressing nine specific emotions [16]. Some studies instructed participants to express specific emotions through touch while recording the touch sounds, from which vibrotactile feedback patterns were derived [17]. The aim was to determine if subjective feelings could be communicated via these simple vibrotactile patterns: the average accuracy across four emotions was 57.74 % by a human decoder. In other work, couples were asked to perform gestures conveying six emotions to one another. The couples’ recognition accuracy was 57 % [31]. However, some researchers point out that direct touch as a mode of communication is influenced by context; research indicates that haptic behaviours by themselves are insufficient for recognizing specific emotions [32]. In one notable study, researchers catalogued 23 distinct types of tactile interactions, including hugging, squeezing, and shaking [30]. The findings from this study revealed that tactile behaviours on their own do not effectively distinguish between different emotions. For instance, the act of stroking was associated with expressions of sadness, as well as love or sympathy. The study also highlighted that factors such as the duration and intensity of the tactile interaction need to be adjusted to better differentiate between emotions.
I-B2 Conveying social gestures
Some researchers have investigated the conveyance of gestures to a mannequin or a robot [33, 16]. In previous studies, an artificial skin was developed to cover the arm of a mannequin, and the Logit-Boost algorithm was used to decode eight social touch gestures with an accuracy of 72 %, compared to human recognition accuracy of 90 % [33]. Additionally, the Social Touch Gesture Challenge 2015 was organized, where the highest accuracy achieved in the competition was 70.9 % for recognizing eight-second-long gestures, covering six touch gestures plus a “no touch” condition [15]. This was further explored in [29], a classification accuracy of 72.8 % was achieved in 2.67 seconds. Furthermore, the Haptic Creature [16] was able to distinguish between four types of touch gestures—“pat”, “Stroke”, “Poke”, and “Squeeze”—with an average accuracy of 77 %.
Earlier research identified specific tactile behaviours associated with each emotion through extensive behavioural coding [4]. For instance, sympathy was linked to stroking and patting, anger to hitting and squeezing, disgust to a pushing motion, and fear to trembling [30]. Additionally, previous studies have explored how individuals use tactile hand gestures to communicate emotions, providing evidence that supports the connection between emotions and gesture features such as intensity, temporal frequency, and spatial frequency [34]. This research aimed to investigate the usability of tactile hand gestures for emotional online communication.
I-C Research goal
Previous studies primarily focused on either tactile or auditory channels in isolation and examined the correlations between social touch gestures and emotions. In this work, we explore the consistency of expressions of emotion based on touch and sound features using multimodal methods. We also concentrate on decoding different emotions without providing specific social gesture options, allowing participants to express different emotions without being restricted to certain gestures.
We focus on the following research questions (RQs):
1.
Can we create touch-sensing capabilities for the Pepper robot?
2.
Are different people consistent in the way they express emotions and gestures through touch?
3.
Which emotions and gestures are distinguishable based on touch? Which ones are easier to decode?
4.
Can multimodal models be used to decode different emotions and gestures? Is a multimodal model better than a unimodal model?
Section II outlines the experimental design for human-robot interaction, including research materials (RQ1), experimental procedures, as well as data collection and processing methods. In Section III, the collected data is analysed. We examine to what extent emotions and gestures are expressed consistently by multiple participants (RQ2), as well as the participants’ subjective notions of correlations between emotions and gestures. Afterwards, Section IV presents the decoding of emotions and gestures using unimodal and multimodal data-driven methods (RQ3, RQ4). Section V discusses our findings, reflects on the limitations of our research, and the paper concludes with Section VI.
II Materials & Methods
This study explores the consistency of touch-based expression of emotions and touch gestures toward the robot by integrating tactile and auditory signals.
II-A Hardware materials
@@
In this experiment, participants were tasked with conveying emotions to a Pepper robot, see Fig. 1. The participants were presented with definitions of the emotions to ensure they agreed with and understood the given definitions. The data collection involved the use of a custom 5-by-5 tactile sensor and a microphone to capture gesture data. Fig. 1(a) shows how the sensors are mounted to the Pepper robot. The following sections will detail the equipment used, the setup process, and the specific configurations implemented for data acquisition.
II-A1 Tactile sensor
@@
The tactile sensor is a 5-by-5 piezoresistive pressure grid, similar to the Smart Textile developed in [35]. As shown in Fig 2(a), the sensor consists of a top and bottom electrode pattern with a piezoresistive Velostat™ sheet in between. The Velostat™ sheet was roughened by pressing it between two pieces of 80 grit sanding paper: this ensures little contact between the Velostat™ and the electrodes when no pressure is applied, and thus a small nominal sensor reading. The electrodes are made by screenprinting DuPont™ PE874™ conductive ink onto a Bemis™ 3914 thermoplastic polyurethane (TPU) sheet. For each electrode, the printed side is facing the Velostat™ sheet. To interface the printed TPU with a printed circuit board (PCB) for readout, flexible PCBs are glued to the printed TPU sheet. This is achieved by adding extra conductive ink on the traces and Gorilla® Super Glue in between. Finally, extra TPU is used to cover the remaining exposed traces, to prevent oxidisation. This was omitted from Fig. 2 for clarity of the graphic. The entire structure is fused together using a Hotronix® Air Fusion IQ® heat press at and 1.4 bar for .
@@
The readout PCB, shown in Fig. 2(b), uses a NINA-B306 microcontroller unit (MCU), powered by an external lithium-ion polymer (LiPo) battery. Flexible printed cable (FPC) connectors are used to interface with the flex PCBs. The MCU communicates the sensor data wirelessly over Bluetooth Low Energy (BLE) at an average rate of 45 Hz. Custom Arduino board files were developed for the PCB so that it can be programmed using the Arduino IDE. We also provide a guide for creating such a custom Arduino board in this GitHub repository111https://github.com/RemkoPr/airo-nrf52840-boards.
@@
The transduction mechanism of the sensor is as follows. When pressure is applied to the sensor, the resistance of the Velostat™ sheet decreases. Hence, as shown in Fig. 2(c), the sensor can be modelled as a grid of variable resistors. Each of the traces of the top electrode is connected to a digital pin D0..5 of the MCU. The traces of the bottom electrode run into a demultiplexer (DEMUX), of which the output is connected to an analog pin A0 of the MCU. This pin has a 1.4 k pulldown resistor. Such a low pulldown value biases the sensor as a sensitive touch sensor, rather than a pressure sensor with a large dynamic range. To read a single taxel, one of the digital pins is set high (3.3 V), the others to a high impedance (Hi-Z) state. The DEMUX then routes one of the bottom traces to the A0 pin. The voltage at this pin depends on the pulldown resistor and the resistance of the Velostat™ squeezed between the top trace set to a voltage level HIGH, and the bottom trace selected by the DEMUX:
II-A2 Microphone
@@
II-B Experimental design
@@
To explore emotion decoding via touch-based expression and reveal how consistent emotions and gestures are expressed across different individuals, we set up a data collection experiment, which can be seen in Fig. 1. We recruited all the participants from a similar cultural background. Specifically, the group comprises 28 Chinese participants, 10 identified as female, 18 as male, and their average age is years. The data collection and study adhered to the ethics procedures of Ghent University, and participants gave informed consent.
II-B1 Definitions of emotions
Affective states can be represented along dimensions of arousal and valence, as described by Russell’s model [36]. In our study, we selected nine distinct emotions from different arousal and valence zones, plus one neutral emotion, as illustrated in Fig. 3, “Anger”, “Fear”, and “Disgust” fall within the high arousal and negative valence quadrant (Quadrant 2). “Happiness” and “Surprise” are in the high arousal and positive valence quadrant (Quadrant 1). “Sadness” and “Confusion” belong to the low arousal and negative valence quadrant (Quadrant 3), while “Comfort” and “calming” are from the low arousal and positive valence quadrant (Quadrant 4). “Attention” is categorized as a neutral emotion (Quadrant 0). Before data collection, participants were provided with definitions of these ten emotions to ensure they understood and agreed with the interpretations or could offer their own understandings. Details of each emotion definition refer to the Oxford dictionary and can be found at this GitHub repository222https://github.com/qiaoqiao2323/Convey-emotion.
II-B2 Data collection procedure
The experiment followed a within-subject design using two sessions: A) Conveying emotions to the robot, and B) Conveying tactile gestures to the robot.
Conveying emotions to the robot: In this session, participants were asked to convey each of the 10 emotions to the robot, one at a time, using spontaneous gestures, with each gesture lasting 10 seconds.
Conveying tactile gestures to the robot: In this session, participants were instructed to convey each of the six gestures to the robot, one at a time, with each gesture lasting 10 seconds.
Participants were given time between each emotion and gesture to prepare and think about their next performance. Subsequent data collection only commenced once the participants indicated they were ready. Participants completed three rounds, with each subsequent round commencing only after all emotions or gestures had been performed, ensuring a greater variety in the data collected. To prevent bias or influence on the gestures used by participants to convey emotions, session A was consistently conducted prior to session B.
II-B3 Questionnaire
After collecting gesture data, we gathered subjective feedback through a questionnaire aimed at understanding the challenges of affective touch expression and the correlation between affective touch and tactile gestures:
Challenging emotions to express: We asked participants which emotions or intentions they find more challenging to express through a single multiple-choice question: “Which emotions or intentions do you find most difficult to express?”
Similar emotions based on touch expression: We gathered pairs of similar emotions by asking participants: “Are there any emotions listed below that you think are similar, based solely on how you express them? Please list your answers in pairs, such as Emotion A – Emotion B.”
Correlations between emotions and gestures: We examined the correlations between emotions and gestures by asking: “Do you believe that certain gestures are associated with specific emotions? If so, which gestures correspond to which emotions?”
General feedback: Finally, we pose an open question: “Do you have suggestions or feedback for this experiment?”
II-C Data processing
Data processing and analysis were done using Python and R V4.3.2. Audio and tactile features were extracted using Python and then analyzed using R. A Permutational Multivariate Analysis of Variance (PERMANOVA) analysis was used to compare data between conditions. A Shapiro-Wilk analysis (with ) was used to confirm the normality of distributions.
II-C1 Dataset
@@
The dataset consists of tactile sensor and microphone recordings collected to study the conveyance of emotions through touch and audio. For each of the 10 emotions, there are a total of 84 samples (28 participants for 3 rounds). Each interaction lasted 10 seconds. Tactile sensor data were sampled at a rate of 45 Hz, providing continuous feedback during the interactions. Simultaneously, 44.1 kHz audio data were recorded in 10-second WAV files.
II-C2 Audio Features
@@ II-C3 Tactile Features
Median Force, representing the middle value of force applied, IQR Force measures the dispersion of the middle 50% of force values. Contact Area is crucial in identifying the nature of the touch, with larger areas potentially indicating comforting emotions, while smaller areas might suggest precision or nervousness. The Rate of Pressure Change captures how quickly the pressure varies, with high rates linked to emotions like excitement. Pressure Std reflects the fluctuation in pressure, and Num_touches quantifies the frequency of touch events within 10 seconds. Max Touch Duration, Min Touch Duration, and Mean Touch Duration describe the temporal aspects of touch, indicating how long touches are sustained, which can be correlated with different emotional expressions.
III Data analysis
III-A Subjective Feedback Analysis on Touch-based Emotional Expression
In our experiment, participants provided subjective feedback on emotional expression through touch. These areas included challenging emotions for expression, similarities in emotional expression based on touch, and correlations between emotions and gestures.
III-A1 Challenging Emotions for Expression
Participants found certain emotions more challenging to express through touch. The results, as shown in Fig. 4(a), indicate that “Surprise” and “Confusion” were the most difficult emotions to convey. Specifically, 64.3% of participants identified “Confusion” as challenging, while 46.4% reported difficulty expressing “Surprise”. These emotions stood out among the ten evaluated in the study. On the contrary, no participants regarded “Attention” as challenging to convey.
III-A2 Similar Emotions Based on Touch Expression
When asked to identify emotions that felt similar based on touch expression alone, participants highlighted several pairs. Fig. 4(b) shows that 14 participants chose “Calming” and “Comfort” as similar emotions. These emotions were consistently associated with low-frequency touch movements and light force. Additionally, 9 participants found “Disgust” and “Anger” to be similar, likely due to both emotions being characterized by high arousal and negative valence. We found that all the similar affective touch pairs given by the participants always belong to the same arousal or sample valence zone; the lack of social context in the experiment may have contributed to the difficulty in differentiating these emotions through touch alone.
III-A3 Correlations Between Emotions and Gestures
Participants’ feedback revealed interesting correlations between emotions and specific gestures. As illustrated in Fig. 4(c), each emotion could be linked to multiple gestures, suggesting a flexible strategy for expressing emotions. For instance, 16 participants chose holding as a gesture for “Calming”, while 7 participants preferred “Rubbing”. Surprisingly, 21 participants associated tapping with “Attention”, and 11 participants chose “Poking”. “Confusion” was linked to various gestures, but none were strongly correlated. On the other hand, “Attention”, “Comfort”, and “Calming” showed stronger associations with specific gestures.
III-A4 Suggestions and feedback for the experiment
Participants provided valuable suggestions for improving the experiment. Some recommend increasing the touch area to enhance emotional expression. Feedback also indicated that considering social roles and context, such as gender, relationship, and appearance, could influence how emotions are conveyed. Participants suggested varying the touch positions and incorporating external stimuli for a more comprehensive analysis. Specific gestures were also highlighted: two participants mentioned using “Stroking” for “Comfort” and “Sadness”, six indicated hitting to express “Anger”, and two suggested pushing. Additionally, several participants expressed concerns about the force used on the robot when expressing “Anger” and “Disgust”, fearing potential damage, and requested guidelines on the safe amount of force to apply during the experiment.
The feedback highlights the complexity of expressing emotions through touch. The varying difficulty in conveying different emotions and the identification of similar arousal and valence zone emotions based on touch alone is challenging. The correlations between emotions and gestures suggest that participants use a combination of gesture features, such as amplitude, frequency, and contact area, to convey different emotions effectively.
III-B Consistency analysis